OpenStudy (anonymous):

The midpoint of line UV is (5, –11). The coordinates of one endpoint are U(3, 5). Find the coordinates of endpoint V.

OpenStudy (anonymous):

\[((3+x_{v})/2, (5 + y _{v})/2) = (5, -11)\] therfore the x-value of v is 7 (as 3+7/2 = 5) and then the y-value of v is -27 (as 5 - 27/2 = -11) so V is (7,-27)